# Learning Mixtures of Linear Dynamical Systems (MoLDS) via Hybrid Tensor–EM Method

### v1

@@ 6 Conclusion
MoLDS offers an interpretable way to model repeated, trajectory-level dynamics in heterogeneous neural recordings. Here, we propose a Tensor-EM algorithm for inference of MoLDS; we use SMD for tensor decomposition as an initialization with Kalman filter-smoother EM, to achieve reliable parameter recovery and interpretable clustering of synthetic and real neural recordings. This hybrid framework makes MoLDS practically usable for large, noisy datasets. As limitations, we note that we have only explored synthetic and real-world cases where the data is likely to have linear dynamics - the synthetic data is linear and the neural data consists of smoothed firing rates - and have not explored cases with high likelihood of model mismatch. Regardless, this study sets the stage for broader applications of MoLDS to complex systems.
@@ 8 Reproducibility Statement
Our work can be reproduced in a straightforward way. The datasets are provided in Miller (2022) and Perich et al. (2018), with all pre-processing techniques detailed in Chowdhury et al. (2020), Lawlor et al. (2018a), and the Appendix. Moreover, detailed algorithms and technical details are provided for each step of the inference, with comprehensive pseudo-code for the implementation in the main text and Appendix.
Appendix A Tensor Initialization Details
In this section, we provide a brief introduction to MoLDS’s connection to MLR and the tensor-based moment method for MoLDS (see Bakshi et al. (2023); Rui & Dahleh (2025) for more details).
A.1 MoLDS to MLR reformulation
The key insight enabling tensor methods for MoLDS is that any linear dynamical system can be equivalently represented through its impulse response (also called Markov parameters), which describe how the system responds to unit impulses in its inputs. This representation allows us to reformulate MoLDS as MLR, making algebraic moment-based methods applicable.
@@ A.1 MoLDS to MLR reformulation
$g_{k}(j)\;=\;C_{k}A_{k}^{\,j-1}B_{k}\;\in\;\mathbb{R}^{m},\qquad j=1,2,3,\ldots,$ (7)
@@
$y_{t}=\sum_{j=1}^{\infty}g_{k}(j)\,u_{t-j}+\text{noise terms}.$ (8)
@@
$y_{i,t}\approx\langle g_{k_{i}}^{(L)},\bar{u}_{i,t}\rangle+\text{noise terms},$ (11)
@@
Step 3: Normalization and MLR formulation.
@@
$\tilde{y}_{j}\;=\;\langle v_{j},\beta_{k_{j}}\rangle+\eta_{j}+\xi_{j},$ (12)
@@
Step 4: Mixture structure.
@@
Thus the learning task reduces to recovering the mixture weights $\{p_{k}\}$ and regression vectors $\{\beta_{k}\}$ from the dataset $\{(v_{j},\tilde{y}_{j})\}$. Once these regression parameters are estimated, the corresponding state-space models $(A_{k},B_{k},C_{k})$ can be reconstructed via the Ho–Kalman realization algorithm Oymak & Ozay (2019).
This reformulation is crucial: it transforms the original problem of identifying a mixture of LDSs into the algebraic problem of learning an MLR model, for which polynomial-time tensor methods with identifiability and sample-complexity guarantees are available Rui & Dahleh (2025).
A.2 Moment construction and whitening
Let $d=Lm$ denote the dimension of the lagged input vectors. We partition the sample indices into two disjoint subsets $\mathcal{N}_{2}$ and $\mathcal{N}_{3}$ for constructing second- and third-order moments, respectively. The empirical moments are defined as
@@ A.2 Moment construction and whitening
$\mathbb{E}[M_{2}]\;=\;\sum_{k=1}^{K}p_{k}\,\beta_{k}\beta_{k}^{\top},\qquad\mathbb{E}[M_{3}]\;=\;\sum_{k=1}^{K}p_{k}\,\beta_{k}^{\otimes 3},$ (14)
@@
$\widehat{T}\;=\;M_{3}(W,W,W)\;=\;\sum_{k=1}^{K}p_{k}\,\alpha_{k}^{\otimes 3},\qquad\alpha_{k}:=W^{\top}\beta_{k}.$ (16)
This tensor is symmetric and orthogonally decomposable. We then apply Simultaneous Matrix Diagonalization (SMD) (A.3) to recover $\{\alpha_{k},p_{k}\}$. Finally, we unwhiten to obtain $\beta_{k}=W^{-\top}\alpha_{k}$, and recover the state-space parameters $(A_{k},B_{k},C_{k})$ via Ho–Kalman realization Oymak & Ozay (2019).
A.3 Simultaneous Matrix Diagonalization
This section provides the Simultaneous Matrix Diagonalization (SMD) method for tensor decomposition Kuleshov et al. (2015) in Algorithm 2. SMD recovers tensor components by reducing the problem to joint matrix diagonalization, exploiting linear-algebraic structure to recover all components simultaneously rather than sequentially as in RTPM.
@@ A.3 Simultaneous Matrix Diagonalization
A.4 Tensor Initialization Algorithm
With the MLR reformulation and SMD tensor method, the full algorithm of tensor initialization for MoLDS is provided below.
@@ A.4 Tensor Initialization Algorithm
22:  Build Hankel matrix from $\{\widehat{g}_{k}^{(h)}\}_{h=1}^{L}$ and apply Ho–Kalman algorithm
@@
Under standard identifiability and excitation conditions, RTPM-based tensor decomposition provably recovers the parameters $\{\beta_{k}\}$ with finite-sample guarantees Rui & Dahleh (2025). Simultaneous matrix diagonalization (SMD) enjoys analogous guarantees in the orthogonal-tensor setting and, in practice, tends to be more numerically stable and noise-robust Kuleshov et al. (2015). These advantages lead to improved parameter estimates during the tensor initialization stage (Figure 1, main paper), which in turn can provide a stronger starting point for the subsequent EM refinement.
Appendix B EM for MoLDS: Complete Technical Details
B.1 Overview and Structure
@@ B.2 Complete EM Algorithm
where the closed-form LDS parameter updates are derived in B.3.
@@
17:   Update LDS parameters $(\hat{A}_{k},\hat{B}_{k},\hat{C}_{k},\hat{D}_{k},\hat{Q}_{k},\hat{R}_{k})$ from $S_{k}$ (see B.3)
@@
B.3 Closed-form M-step Updates
The M-step updates follow standard LDS maximum likelihood estimation Ghahramani & Hinton (1996). For simplicity, we present the case $D_{k}=0$ (no direct feedthrough) and drop component and iteration indices $(k,t)$ for readability:
@@ B.3 Closed-form M-step Updates
$R$ $=\frac{1}{T}\big(S_{yy}-CS_{yx}^{\top}-S_{yx}C^{\top}+CS_{xx}C^{\top}\big),$
$Q$ $=\frac{1}{N}\Big(S_{xx,\mathrm{curr}}-AS_{xx^{-}}^{\top}-BS_{xu^{-}}^{\top}-(AS_{xx^{-}}^{\top}+BS_{xu^{-}}^{\top})^{\top}$
$\qquad\qquad+AS_{x^{-}x^{-}}A^{\top}+AS_{u^{-}x^{-}}^{\top}B^{\top}+BS_{u^{-}x^{-}}A^{\top}+BS_{u^{-}u^{-}}B^{\top}\Big).$
@@ B.4 Computational Complexity
Tensor initialization significantly reduces the iteration count compared to random initialization, which substantially improves computational efficiency. The tensor initialization phase requires $O(d^{3})$ operations where $d=Lm$ is the MLR dimension, making it negligible in the whole Tensor–EM pipeline.
Appendix C Initializing $(Q,R)$ After Tensor Initialization
@@
where we use $\lambda=10^{-6}\max(\mathrm{diag}(\hat{C}_{z}^{\top}\hat{C}_{z}))$ for numerical stability.
@@
With $T_{i}^{\mathrm{tr}}=T_{i}-1$, we set
$\hat{Q}_{z}^{(0)}\;=\;\frac{\sum_{i}w_{i,z}\sum_{t=0}^{T_{i}-2}\eta_{i,t}^{(z)}\eta_{i,t}^{(z)\!\top}}{\sum_{i}w_{i,z}\,T_{i}^{\mathrm{tr}}},\qquad\hat{R}_{z}^{(0)}\;=\;\frac{\sum_{i}w_{i,z}\sum_{t=0}^{T_{i}-1}\varepsilon_{i,t}^{(z)}\varepsilon_{i,t}^{(z)\!\top}}{\sum_{i}w_{i,z}\,T_{i}}.$ (26)
@@
Appendix D Tensor-EM MoLDS Application on Neural Dataset Details
D.1 Area2 Dataset
@@
We evaluate our methods on a neural population dataset recorded from the motor cortex (Area2) of a rhesus macaque. The task is a center-out reaching paradigm where the subject applies force to a manipulandum in response to cues. Neural activity was recorded with a Utah array and spike counts were extracted from thresholded multiunit activity. The dataset provides simultaneously measured behavioral covariates (applied force, hand kinematics) alongside the neural recordings.
@@
Following the Neural Latents Benchmark preprocessing, we first apply principal component analysis (PCA) to reduce the neural activity to $p$ latent dimensions (here $p=6$). The behavioral force signals (6D) are used as exogenous inputs. For MoLDS training, we constructed input-output trials of the form $(u_{t},y_{t})$, where $u_{t}\in\mathbb{R}^{6}$ corresponds to force inputs and $y_{t}\in\mathbb{R}^{p}$ are PCA-reduced neural outputs. Each reaching direction corresponds to a separate trial. We split the dataset into non-overlapping train, val, and test sets, ensuring balanced coverage across each direction.
@@
For model initialization, we computed input-output moment tensors from the training data and applied Simultaneous Matrix Diagonalization (SMD) to obtain globally consistent estimates of system parameters and mixture weights. These estimates were used to initialize a full Kalman filter-smoother EM refinement, which updates all LDS parameters in closed form. We trained MoLDS models with different numbers of mixture components $K\in{3,4,5}$, latent dimension $n\in{3,4,5}$, and lag parameters $L\in{16,24}$. Validation data is used for model selection by comparing negative log-likelihood (NLL), root mean squared error (RMSE), and Bayesian Information Criterion (BIC).
@@
Supervised per-direction baseline.
@@
We fit switching LDS (SLDS) models with $S\in{1,2,3}$ discrete states to the training set (Gaussian emissions, 6-D force inputs, sticky transitions). Validation NLL/BIC selected $S=2$ SLDS. We then decoded the test set with Viterbi and summarized the inferred state sequences (Fig. 6). Panel (a) shows that within-trial switches are infrequent and tend to occur in similar temporal regions across trials. In panel (b, left), grouping test trials by their dominant state $s^{*}$ reveals that the same state dominates most directions, with only a mild departure around dir-$90$. The mean time-share view (b, right) tells the same story: state composition looks broadly similar across directions. Consistent with this, a “collapsed” SLDS that uses only each trial’s dominant state achieves nearly the same predictive metrics as the fully segmented model, indicating that within-trial switching adds little in this dataset. These results support our interpretation that the primary structure is between trials (by direction), not within trials.
Figure 6: SLDS result: (a) Inferred switch states for all test trials. (b) Group trials and mean time by states.
D.2 PMd Dataset

### v2

@@ 6 Conclusion
MoLDS provides an interpretable framework for modeling repeated, trajectory-level dynamics in heterogeneous data. In this work, we proposed a hybrid Tensor-EM approach for efficient and reliable MoLDS inference. Specifically, we used SMD for tensor-based initialization, followed by Kalman filter–smoother EM refinement. This combination enables accurate parameter recovery and interpretable trajectory clustering in both synthetic benchmarks and real neural datasets. This hybrid framework makes MoLDS practically usable for large-scale, noisy datasets.
Our empirical evaluation focused on scenarios with linear dynamics: synthetic data generated from linear systems and neural data consisting of smoothed firing rates over selected movement windows. While these experiments validate the effectiveness of the proposed inference procedure, we have not explored cases with a high likelihood of model mismatch or strong nonlinearities. We envision a few extensions that can help broaden the applicability of MoLDS. First, nonlinear dynamics may be approximated by segmenting trajectories into shorter sessions that are locally well-approximated by linear dynamical systems. Second, linear components could be augmented with nonlinear features to capture moderate deviations from linear structure. In addition, future directions can include extending the framework to autonomous settings and developing theoretical guarantees under simplified assumptions.
More broadly, MoLDS is well-suited to various neuroscience tasks in which multiple experimental conditions share latent dynamical structure but differ in observable trajectories. For example, in sensory decision-making with different stimulus strengths, distinct conditions may reuse a small set of evidence-integration dynamics (Hanks et al., 2015). In context-switching tasks, neural populations in the prefrontal cortex may transition among a number of rule-dependent dynamical modes (Mante et al., 2013). In motor learning experiments, early, intermediate, and late learning stages may correspond to transitions among a small set of dynamical regimes (Shenoy et al., 2013). By developing a principled Tensor-EM learning and inference framework, this work establishes a practical and extensible foundation for applying MoLDS to complex neural datasets. Beyond neuroscience, the framework may also support broader applications, such as identifying clinically relevant behavioral dynamics (Bulteel et al., 2016) or uncovering cell-type-specific dynamical structure (Luecke et al., 2021).
@@ 8 Reproducibility Statement
Our work can be reproduced in a straightforward way. The datasets are provided in (Miller, 2022) and (Perich et al., 2018), with all pre-processing techniques detailed in (Chowdhury et al., 2020), (Lawlor et al., 2018), and in the Appendix of this work. Moreover, detailed algorithms and technical details are provided for each step of the learning and inference, with comprehensive pseudo-code for the implementation in the main text and the Appendix.
9 Acknowledgments
L. Gong was supported by the Swartz Foundation Postdoctoral Fellowship. S. Saxena was supported by the National Science Foundation grant No. 2350329 and the National Institutes of Health grant No. R34DA059718. The authors especially thank Maryann Rui and Dr. Munther Dahleh for helpful discussions.
Appendix A LLM Usage Statement
We used LLMs to assist with language editing in parts of the manuscript. All LLM’s suggestions were manually verified and edited by the authors.
Appendix B Tensor Initialization Details
In this section, we provide a brief introduction to MoLDS’s connection to MLR and the tensor-based moment method for MoLDS (see (Bakshi et al., 2023; Rui & Dahleh, 2025) for more details).
B.1 MoLDS to MLR reformulation
The key insight enabling tensor methods for MoLDS is that any linear dynamical system can be equivalently represented through its impulse response (also called Markov parameters), which describes how the system responds to unit impulses in its inputs. This representation allows us to reformulate MoLDS as MLR, making algebraic moment-based methods applicable. We now briefly introduce this process following the formulation in (Rui & Dahleh, 2025).
@@ B.1 MoLDS to MLR reformulation
$g_{k}(j)\;=\;C_{k}A_{k}^{\,j-1}B_{k}\;\in\;\mathbb{R}^{m},\qquad j=1,2,3,\ldots,$
@@
$y_{t}=\sum_{j=1}^{\infty}g_{k}(j)\,u_{t-j}+\text{noise terms}.$
@@
$y_{i,t}\approx\langle g_{k_{i}}^{(L)},\bar{u}_{i,t}\rangle+\text{noise terms},$
@@
Step 4: Normalization and MLR formulation.
@@
$\tilde{y}_{j}\;=\;\langle v_{j},\beta_{k_{j}}\rangle+\eta_{j}+\xi_{j},$
@@
Step 5: Mixture structure.
@@
Thus the learning task reduces to recovering the mixture weights $\{p_{k}\}$ and regression vectors $\{\beta_{k}\}$ from the dataset $\{(v_{j},\tilde{y}_{j})\}$. Once these regression parameters are estimated, the corresponding state-space models $(A_{k},B_{k},C_{k})$ can be reconstructed via the Ho-Kalman realization algorithm (Oymak & Ozay, 2019).
This reformulation is crucial: it transforms the original problem of identifying a mixture of LDSs into the algebraic problem of learning an MLR model, for which polynomial-time tensor methods with identifiability and sample-complexity guarantees are available (Rui & Dahleh, 2025).
B.2 Moment construction and whitening
Let $d=Lm$ denote the dimension of the lagged input vectors. We partition the sample indices into two disjoint subsets $\mathcal{N}_{2}$ and $\mathcal{N}_{3}$ for constructing second- and third-order moments, respectively as in (Rui & Dahleh, 2025). The empirical moments are defined as
$M_{2}\;=\;\frac{1}{2|\mathcal{N}_{2}|}\sum_{j\in\mathcal{N}_{2}}\tilde{y}_{j}^{\,2}\,\big(v_{j}\otimes v_{j}-I_{d}\big),\qquad M_{3}\;=\;\frac{1}{6|\mathcal{N}_{3}|}\sum_{j\in\mathcal{N}_{3}}\tilde{y}_{j}^{\,3}\,\big(v_{j}^{\otimes 3}-\mathcal{E}(v_{j})\big),$
@@ B.2 Moment construction and whitening
$\mathbb{E}[M_{2}]\;=\;\sum_{k=1}^{K}p_{k}\,\beta_{k}\beta_{k}^{\top},\qquad\mathbb{E}[M_{3}]\;=\;\sum_{k=1}^{K}p_{k}\,\beta_{k}^{\otimes 3},$
@@
$\widehat{T}\;=\;M_{3}(W,W,W)\;=\;\sum_{k=1}^{K}p_{k}\,\alpha_{k}^{\otimes 3},\qquad\alpha_{k}:=W^{\top}\beta_{k}.$
This tensor is symmetric and orthogonally decomposable. We then apply Simultaneous Matrix Diagonalization (SMD) (B.3) (Kuleshov et al., 2015) to recover $\{\alpha_{k},p_{k}\}$. Finally, we unwhiten to obtain $\beta_{k}=W^{-\top}\alpha_{k}$, and recover the state-space parameters $(A_{k},B_{k},C_{k})$ via Ho-Kalman realization (Oymak & Ozay, 2019).
B.3 Simultaneous Matrix Diagonalization
This section provides the Simultaneous Matrix Diagonalization (SMD) method for tensor decomposition (Kuleshov et al., 2015) in Algorithm 2. SMD recovers tensor components by reducing the problem to joint matrix diagonalization, exploiting linear algebraic structure to recover all components simultaneously rather than sequentially as in RTPM.
@@ B.3 Simultaneous Matrix Diagonalization
B.4 Tensor Initialization Algorithm
With the MLR reformulation and the SMD tensor method, the full algorithm of tensor initialization for MoLDS can be provided now.
@@ B.4 Tensor Initialization Algorithm
22:  Build Hankel matrix from $\{\widehat{g}_{k}^{(h)}\}_{h=1}^{L}$ and apply Ho-Kalman algorithm
@@
Under standard identifiability and excitation conditions, RTPM-based tensor decomposition provably recovers the parameters $\{\beta_{k}\}$ with finite-sample guarantees (Rui & Dahleh, 2025). Simultaneous matrix diagonalization (SMD) enjoys analogous guarantees in the orthogonal-tensor setting and, in practice, tends to be more numerically stable and noise-robust (Kuleshov et al., 2015). These advantages lead to improved parameter estimates during the tensor initialization stage (Figure 2, main paper), which in turn can provide a stronger starting point for the subsequent EM refinement.
Appendix C EM for MoLDS: Complete Technical Details
C.1 Overview and Structure
@@ C.2 Complete EM Algorithm
where the closed-form LDS parameter updates are derived in C.3.
@@
17:   Update LDS parameters $(\hat{A}_{k},\hat{B}_{k},\hat{C}_{k},\hat{D}_{k},\hat{Q}_{k},\hat{R}_{k})$ from $S_{k}$ (see C.3)
@@
C.3 Closed-form M-step Updates
The M-step updates follow standard LDS maximum likelihood estimation (Ghahramani & Hinton, 1996). For simplicity, we present the case $D_{k}=0$ (no direct feedthrough) and drop component and iteration indices $(k,t)$ for readability:
@@ C.3 Closed-form M-step Updates
$R$ $=\frac{1}{T}\big(S_{yy}-CS_{yx}^{\top}-S_{yx}C^{\top}+CS_{xx}C^{\top}\big),$ (17)
$Q$ $=\frac{1}{N}\Big(S_{xx,\mathrm{curr}}-AS_{xx^{-}}^{\top}-BS_{xu^{-}}^{\top}-(AS_{xx^{-}}^{\top}+BS_{xu^{-}}^{\top})^{\top}$ (18)
$\qquad\qquad+AS_{x^{-}x^{-}}A^{\top}+AS_{u^{-}x^{-}}^{\top}B^{\top}+BS_{u^{-}x^{-}}A^{\top}+BS_{u^{-}u^{-}}B^{\top}\Big).$ (19)
@@ C.4 Computational Complexity
Tensor initialization significantly reduces the iteration count compared to random initialization, which substantially improves computational efficiency. The tensor initialization phase requires $O(d^{3})$ operations where $d=Lm$ is the MLR dimension, making it negligible in the whole Tensor-EM pipeline.
Appendix D Initializing $(Q,R)$ After Tensor Initialization
@@
where we recommend $\lambda=10^{-6}\max(\mathrm{diag}(\hat{C}_{z}^{\top}\hat{C}_{z}))$ for numerical stability.
@@
With $T_{i}^{\mathrm{tr}}=T_{i}-1$, set
$\hat{Q}_{z}^{(0)}\;=\;\frac{\sum_{i}w_{i,z}\sum_{t=0}^{T_{i}-2}\eta_{i,t}^{(z)}\eta_{i,t}^{(z)\!\top}}{\sum_{i}w_{i,z}\,T_{i}^{\mathrm{tr}}},\qquad\hat{R}_{z}^{(0)}\;=\;\frac{\sum_{i}w_{i,z}\sum_{t=0}^{T_{i}-1}\varepsilon_{i,t}^{(z)}\varepsilon_{i,t}^{(z)\!\top}}{\sum_{i}w_{i,z}\,T_{i}}.$
@@
Appendix E Tensor-EM MoLDS Application on Neural Dataset Details
E.1 Area2 Dataset
@@
We evaluate our methods on a neural population dataset recorded from the motor cortex (Area2) of a rhesus macaque. The task is a center-out reaching paradigm where the subject applies force to a manipulandum in response to cues. Neural activity was recorded with a Utah array, and spike counts were extracted from thresholded multiunit activity. The dataset provides simultaneously measured behavioral covariates (applied force, hand kinematics) alongside the neural recordings.
@@
Following the Neural Latents Benchmark preprocessing, we first apply principal component analysis (PCA) to reduce the neural activity to $p$ latent dimensions (here we use $p=6,20$). The behavioral force signals (6D) are used as exogenous inputs. For MoLDS training, we constructed input-output trials of the form $(u_{t},y_{t})$, where $u_{t}\in\mathbb{R}^{6}$ corresponds to force inputs and $y_{t}\in\mathbb{R}^{p}$ are PCA-reduced neural outputs. Each reaching direction corresponds to a separate trial. We split the dataset into non-overlapping train, val, and test sets, ensuring balanced coverage across each direction.
We checked the variance explained by PCA and found that the first six PCs capture over $90\%$ of the total variance for each direction, while the first 20 PCs explain nearly all of it. This indicates that PCA preserves the essential structure of the neural activity. The following figure illustrates this.
Figure 7: Area2 neural activity PCA spectrum.
@@
For model initialization, we computed input-output moment tensors from the training data and applied SMD to obtain globally consistent estimates of system parameters and mixture weights. These estimates were used to initialize a full Kalman filter-smoother EM refinement, which updates all LDS parameters in closed form. We trained MoLDS models with different numbers of mixture components $K\in{3,4,5}$, latent dimension $n\in{3,4,5}$, and lag parameters $L\in{16,24}$. Validation data is used for model selection by comparing negative log-likelihood (NLL), root mean squared error (RMSE), and Bayesian Information Criterion (BIC).
@@
Supervised per-direction LDS fitting baseline.
@@
Single-trial LDS fitting baseline.
In addition, we also fit an LDS for each single trial. We compute the average LDS impulse response in each direction and compare the obtained clusters as compared to the per-direction LDS. Following the same procedure as in per-dir LDS fitting, we compare the clusters of single-trial LDS fitting with MoLDS components as well.
Figure 8: Markov parameters representation of LDS models from different fitting methods.
We have examined the trial-to-trial variability of the Area2 dataset by comparing their full Markov parameter vectors between single-trial, perd-dir, and single global LDS fits. As shown in Figure 8, by evaluating the first three PCs of the Markov vectors, single-trial LDSs within each direction form compact clusters, and the pooled per-direction LDS lies near the cluster center for most directions. In contrast, these single-trial LDS are very far from a global LDS fitted on all trials. This indicates that within-direction variability is mild, and the per-direction LDS provides a stable summary of the underlying dynamics rather than an oversimplification. We also performed clustering directly on independently fitted single-trial LDSs. These LDSs naturally cluster by movement direction, even without pooling, demonstrating that direction-specific dynamical signatures are already present at the single-trial level. Moreover, by comparing with the per-dir LDS and MoLDS fit, these single-trial LDS fit-based clusters are very similar, with only two directions being switched between clusters (see Figure 12). Together, these analyses show that while the per-direction LDS is not a perfect “gold standard,” it is a reasonable and stable reference, and our comparisons to MoLDS capture meaningful structure in the data.
@@
We fit switching LDS (SLDS) models with $S\in{1,2,3}$ discrete states to the training set (Gaussian emissions, 6-D force inputs, mild or strong sticky transitions). Validation NLL/BIC selected $S=2$ SLDS. We then decoded the test set with Viterbi and summarized the inferred state sequences (Figure 9). Panel (a) shows that within-trial switches are infrequent and tend to occur in similar temporal regions across trials. In panel (b, left), grouping test trials by their dominant state $s^{*}$ reveals that the same state dominates most directions, with only a mild departure around dir-$90$. The mean time-share view (b, right) tells the same story: state composition looks broadly similar across directions. Consistent with this, a “collapsed” SLDS that uses only each trial’s dominant state achieves nearly the same predictive metrics as the fully segmented model, indicating that within-trial switching adds little in this dataset. These results support our interpretation that the primary structure is between trials (by direction), not within trials.
Figure 9: SLDS result: Inferred switch states for all test trials with mild (a) and strong stickiness (b). (a) Group trials and mean time by states with the mildly sticky SLDS fit.
More results.
Impulse response comparison.
 We compare the direction-specific impulse response (Markov parameter) curves of the assigned MoLDS component and the corresponding single LDS fit for that direction, which shows high similarity (see Figure 10 where titles report cosine similarity).
Figure 10: Impulse response comparison between MoLDS components and per-dir LDS fit.
Results of MoLDS fitting using 20 PCs.  We also applied the MoLDS method on the Area2 dataset with 20 PCs, and found consistent results for the 6-PC fitting model as shown in Figure 11.
Figure 11: MoLDS application on Area2 with 20 PCs: results are consistent with those presented in the main text.
Multi-step prediction  In addition to the one-step-ahead predictions presented in the main paper, we also inspected multi-step and free-run predictions. A representative example is provided in Figure 12. The model follows the true neural trajectory for several steps before diverging as it runs forward without correction, which is the characteristic and expected pattern for linear dynamical systems. This behavior is consistent with LDS-based modeling and does not affect our use of MoLDS for identifying shared local dynamical structure across trials.
Figure 12: Multi-step forward predictions of MoLDS on one trial of Area2 data: 5-step vs fully free run after 100ms warm-up.
Responsibility distribution. We analyzed the full responsibility distribution of MoLDS components on held-out test trials to assess whether non-dominant components carry substantial probability. Across all trials, the dominant responsibility is very large $0.976$, while the second-largest responsibility is very small $0.023$. This pattern is consistent across directions as shown in Figure 13.
Figure 13: Area2 MoLDS fit dominant and second component analysis.
E.2 PMd Dataset